Jordan Company is considering the addition of a new product to its cosmetics line. The company has three distinctly different options a skin cream, a bath oil, or a hair coloring gel. Relevant information and budgeted annual income statements for each of the products follow Relevant Information Skin Cream 120,000 10 2 Bath Oil 200,000 6 Color Gel 80,900 13 Budgeted sales in units (a) Expected sales price (b) Variable costs per unit (c) Income statements Sales revenue (a x b) Variable costs (a x c) Contribution margin Fixed costs Net income $1,200,000 (240,000) 960,000 $1,200,000 $1,040,000 (600,000) (640,000) 600,000 (609,000)435,900 $ 165,000 400,000 120,000 $ 280, 000 $ 351,000 Required: a. Determine the margin of safety as a percentage for each product. b. Prepare revised income statements for each product, assuming a 20 percent increase in the budgeted sales volume c. For each product, determine the percentage change in net income that results from the 20 percent increase in sales d. Assuming that management is pessimistic and risk averse, which product should the company add to its cosmetics line? e. Assuming that management is optimistic and risk aggressive, which product should the company add to its cosmetics line?

Req A Req B Req C Req D to E Determine the margin of safety as a percentage for each product. (Round your answers to whole percentage values.) Skin Cream Bath Oil Color Gel Margin of safety

Req A Req B Req C Req D to E Prepare revised income statements for each product, assuming a 20 percent increase in the budgeted sales volume JORDAN COMPANY Income Statements Skin Cream Bath Oil Color Gel Sales revenue Variable costs Contribution margin Fixed cost Net income 0 S

Req A Req B Req C Req D to E For each product, determine the percentage change in net income that results from the 20 percent increase in sales. (Round your answers to whole percentage values.) Skin Cream Bath Oil Color Gel Percentage change in net income

Assuming that management is pessimistic and risk averse, which product should the company add to its cosmetics line? Assuming that management is optimistic and risk aggressive, which product should the company add to its cosmetics line?
Answer #1

a

Calculation of margin of safety percentage

For skin cream

Contribution margin ratio = Contribution margin/Sales

= 960,000/1.200,000

= 80%

Margin of safety = Profit/Contribution margin ratio

= 351,000/80%

= $438,750

Margin of safety (%) = Margin of safety/Sales

= 438,750/1,200,000

= 36.56%

For bath oil

Contribution margin ratio = Contribution margin/Sales

= 600,000/1,200,000

= 50%

Margin of safety = Profit/Contribution margin ratio

= 165,000/50%

= $330,000

Margin of safety (%) = Margin of safety/Sales

= 330,000/1,200,000

= 27.5%

For color gel

Contribution margin ratio = Contribution margin/Sales

= 400,000/1,040,000

= 38.46%

Margin of safety = Profit/Contribution margin ratio

= 280,000/38.46%

= $728,030

Margin of safety (%) = Margin of safety/Sales

= 728,030/1,040,000

= 70%

(b) Revised income statement after 20% increase in sales

Due to 20% increase in sales, variable cost will also increase by 20%.

Skin cream

Bath oil

Color gel

Sales

1,440,000

1,440,000

1,248,000

Less: Variable cost

- 288,000

- 720,000

- 768,000

Contribution margin

1,152,000

720,000

480,000

Less: Fixed cost

- 609,000

- 435,000

- 120,000

Profit

543,000

285,000

360,000

(c)

Calculation of percentage increase in net income

Percentage increase in net income of skin cream = (543,000 - 351,000)/351,000

= 54.70 %

Percentage increase in net income of bath oil = (285,000 - 165,000)/165,000

= 72.73 %

Percentage increase in net income of color gel = (360,000 - 280,000)/280,000

= 28.57%

Calculation of operating leverage

Operating leverage = Contribution/Net income

Operating leverage for skin cream = 960,000/351,000

= 2.74

Operating leverage for bath oil = 600,000/165,000

= 3.64

Operating leverage for color gel = 400,000/280,000

= 1.43

(d) If the management is pessimistic and risk averse, company will add color gel to its cosmetic line since operating leverage is least in case of color gel. Lower operating leverage implies lower risk and low income potential.

(f) If the management is optimistic and risk aggressive, company will add bath oil to its cosmetic line since operating leverage is highest in case of bath oil. Higher operating leverage implies higher risk and higher income potential

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.